# Minimax rates in variance and covariance changepoint testing

### v1

@@ 4.1 A convex relaxation of the k-sparse eigenvalue problem
The testing procedures presented in the previous section have provably strong theoretical performance, but are computationally intractable for all but small values of $p$. Indeed, a key ingredient in these tests is the $s$-sparse eigenvalue of $\widehat{\Sigma}_{1,t}-\widehat{\Sigma}_{2,t}$, which has dimension $p\times p$. Seeing as the sparse eigenvalue is NP hard to compute in general (Magdon-Ismail, 2017), the changepoint testing procedures in the previous section are thus prohibitively computationally costly unless $p$ is very small.
As a remedy to the large computational cost, the $s$-sparse eigenvalue can be approximated by a convex relaxation of the implicit optimization problem, as is done in Berthet and Rigollet (2013) for testing for a rank one perturbation in an isotropic covariance matrix. A similar approach is also taken in Wang and Samworth (2017), where a sparse eigenvector is of interest, but is costly to compute. Following Berthet and Rigollet (2013) and d’Aspremont et al. (2007), we can reformulate the sparse eigenvalue problem as follows. For any symmetric matrix $A\in\mathbb{R}^{p\times p}$, recalling that
@@
The optimization problem in (53) features two sources of non-convexity; the $\ell_{0}$ constraint and the rank constraint. To relax the $\ell_{0}$ constraint, notice that $\left\lVert Z\right\rVert_{1}\leq s$ whenever $\mathrm{Tr}(Z)=1$ and $\left\lVert Z\right\rVert_{0}\leq s^{2}$, since $\left\lVert Z\right\rVert_{1}\leq s\left\lVert Z\right\rVert_{2}=s\mathrm{Tr}(Z)^{1/2}=s\mathrm{Tr}(v^{\top}v)^{1/2}=s$. Thus, a convex relaxation of the $\ell_{0}$ constraint is given by $\left\lVert Z\right\rVert_{1}\leq s$. For the rank constraint, however, there is no obvious convex relaxation. By simply dropping the rank constraint, a convex relaxation of the optimization problem in (53) is given by
@@
where $N(p,s)=\{Z\in\mathbb{R}^{p\times p}\ ;\ Z\succcurlyeq 0,\mathrm{Tr}(Z)=1,\left\lVert Z\right\rVert_{1}\leq s\}$. As a convex problem, it can be solved efficiently using e.g. interior point methods (see Boyd and Vandenberghe, 2004) or first order methods (see Bach et al., 2010), the latter in polynomial time in $p$.
Due to the results in Bach et al. (2010), formalized in Lemma 11 in the Supplementary Material, we have
@@ 4.2 The testing procedure
where the $\widehat{\Sigma}_{i,t}$ for $i=1,2$ are defined in (41). Due to slightly larger deviation bounds on $\widehat{S}_{t,s}$ under the null, one can only guarantee detection of a changepoint whose location $t_{0}$ satisfies $\min(t_{0},n-t_{0})\geq\beta$, where $\beta=\beta(p,n,s)=s^{2}\{\log(ep)\vee\log\log(8n)\}$. Hence, a natural estimate of the noise level using the ”convexified” sparse eigenvalue is given by
$\widehat{\sigma}^{2}_{s,\mathrm{con}}=\widehat{\lambda}^{s}_{\max}(\widehat{\Sigma}_{1,\lceil\beta\rceil})\wedge\widehat{\lambda}^{s}_{\max}(\widehat{\Sigma}_{2,\lceil\beta\rceil}).$ (58)
Here, the first and last $\lceil\beta\rceil$ data points are used to estimate the noise level of the data, as no changepoints in these segments can be guaranteed to be detected by our testing procedure anyway.
Our modified adaptive test is defined by
@@
for some $s\in[p]$, where $C>0$ is some constant depending only on $\lambda,\delta,w$ and $u$.
The signal strength condition (62) in Proposition 6 is larger than that in Proposition 5 by a factor of at least $s$, ignoring absolute constants. Thus, the computationally feasible testing procedure in (59) is near-optimal when the sparsity $s$ is small. Notice, however, that the testing procedure in (48) is computationally feasible when the set of permitted sparsities in (49) is reduced to $\mathcal{S}=\{p\}$, as the sparse eigenvalue reduces to the operator norm in this case. The test in (59) can therefore be augmented with the test in (48), still maintaining low computational cost. Then, Proposition 6 still holds (possibly with different absolute constants) when the right hand side of (62) is replaced by $p\vee\log\log(8n)$ for $s=p$, making the combined testing procedure near-optimal also when $s$ is close to $p$.
Acknowledgments
@@ 5.1 Proofs of results from Section 2
where we in the second inequality used that $\log\log(8n)\geq 1$, and last inequality used that $c_{1}C^{1/2}/(2c_{3})-\lambda>0$ and $t\leq(16c_{0}^{2}\vee 1)\log(8/\delta)$. Due to the definition of $C$, the right hand side of (89) is positive, and therefore $\psi_{\lambda}(X)=1$ holds on the event $\mathcal{E}$ also when $t\leq(16c_{0}^{2}\vee 1)\log(8/\delta)$.
@@
Fix any $\sigma>0,w\geq(2\pi)^{-1/2}$ and $u>0$ sufficiently large. To prove Proposition 2 we will impose a prior on the alternative hypothesis parameter space and bound the total variation distance between the null distribution and the mixture distribution induced by the prior. Since $w\geq(2\pi)^{-1/2}$ and $u$ is sufficiently large, we can take the distribution of the $X_{i}$ to be Gaussian. Given any $0\leq\eta\leq 1$, it then suffices by Lemma 13 to find an absolute constant $c>0$ and a probability measure $\nu$ with $\text{supp}(\nu)\subseteq\Theta(n,c\log\log(8n))$ such that $\chi^{2}(f_{1},f_{0})\leq\eta,$ where $f_{0}$ and $f_{1}$ are the densities of $X\sim\text{N}_{n}(0,V)$ when $V=\sigma^{2}I$ and $V\sim\nu$, respectively.
@@
    Let $\Delta=2^{l}$, where $l$ is sampled uniformly from the set $\left\{0,1,2,\ldots,{\lfloor\log_{2}(n/2)\rfloor\}}\right\}$.
@@
$\min(\Delta,n-\Delta)\left(\frac{\left|\sigma_{1}^{2}-\sigma_{2}^{2}\right|}{\sigma_{1}^{2}\wedge\sigma^{2}_{2}}\right)\wedge\left(\frac{\left|\sigma_{1}^{2}-\sigma_{2}^{2}\right|}{\sigma_{1}^{2}\wedge\sigma^{2}_{2}}\right)^{2}=\rho,$ (92)
@@
where $l_{i}=\log_{2}(\Delta_{i})$ for $i=1,2$. We bound the two terms on the right hand side of (98) separately. For the first term, we have
@@
Now choose $c\leq\log(16/\eta)^{-1}/2\wedge 1$. Then if $\log\log(8n)<c^{-1}$, the right hand side of (100) is zero since $c\log\log(8n)<1$. If we instead have $\log\log(8n)\geq c^{-1}$, the right hand size of (100) is bounded above by
@@
Now we bound the second term in the right hand side of (98). We have
@@
Note that $l_{1},l_{2}$ are sampled independently and uniformly from an integer interval with cardinality $a(n)=\left\lfloor\log_{2}(n/2)\right\rfloor+1$, from which it follows that $\mathbb{P}(|l_{1}-l_{2}|=x)\leq 2a(n)^{-1}$ for any $x\in\mathbb{N}\cup\{0\}$. We claim that the first term at the right hand side of (106) is bounded above by
@@
$\mathbb{E}_{l_{1},l_{2}}\exp\left\{\frac{c}{2}\log\log(8n)\right\}\mathbbm{1}\left\{\left|l_{1}-l_{2}\right|=0\right\}$ $=\exp\left\{\frac{c}{2}\log\log(8n)\right\}\mathbb{P}_{l_{1},l_{2}}(|l_{1}-l_{2}|=0)$ (114)
@@
For the second term at the right hand side of (106), using that $c\leq 1$, we have
@@
Now let $b_{\eta}=\ \underset{n\geq 2}{\sup}\ \frac{\log\log(8n)}{\log(8n)^{\eta\log(2)/18}}<\infty$. Then for the second term on the right hand side of (106), we have
@@ 5.2 Proofs of results from Section 3
    Let $\Delta=2^{l}$, where $l$ is sampled uniformly from the set $\left\{0,1,2,\ldots,\lfloor\log_{2}(n/2)\rfloor\right\}.$
@@
$\min(t_{0},n-t_{0})\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)\wedge\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)^{2}=\rho,$ (126)
@@
Using similar arguments as above to bound the two terms at right hand side of Equation (131), the event
@@
$\mathcal{E}_{5}$ $=\bigcap_{s\in\mathcal{S}}\left\{c_{4}\leq\frac{\widehat{\sigma}^{2}_{s}}{\lambda_{\max}^{s}({\Sigma})}\right\},$ (150)
@@
Since $C\geq 32$, the chain of inequalites preceding Equation (165) imply that $t\geq s\log(ep/s)\vee\log\log(8n)$, since $\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})\leq\lambda_{\max}^{s}(\Sigma_{1})\vee\lambda_{\max}^{s}(\Sigma_{2})$, due to Lemma 12. It follows that
@@ 5.3 Proofs of results from Section 4
$\mathcal{E}_{9}$ $=\bigcap_{s\in\mathcal{S}}\left\{c_{4}\leq\frac{\widehat{\sigma}^{2}_{s,\mathrm{con}}}{\lambda_{\max}^{s}({\Sigma})}\right\},$ (173)
$\mathcal{E}_{10}$ $=\bigcap_{t\in\mathcal{T}}\bigcap_{s\in\mathcal{S}}\left\{\widehat{\lambda}_{\max}^{s}({\widehat{\Sigma}_{1,t}-\widehat{\Sigma}_{2,t}})\leq c_{6}\lambda_{\max}^{s}({\Sigma})h(p,n,s,t)\right\},$ (174)
where $\widehat{\sigma}^{2}_{s,\mathrm{con}}$ is given in (58), $\widehat{\Sigma}_{i,t}$ is given in (41), $\mathcal{T}$ is given in (21), $\mathcal{S}$ is given in (49), $h(p,n,s,t)$ is given in (60), $c_{4}=\delta^{2}(\delta^{2}+16\delta+64)^{-1}(2e\pi w^{2})^{-1}$, $c_{6}=c_{0}\{1+\log(4/\delta)\}$, and $c_{0}$ is the constant from Lemma 2 depending only on $u>0$. Define the event $\mathcal{E}=\mathcal{E}_{9}\cap\mathcal{E}_{10}$. By Lemma 7 we have $\mathbb{P}(\mathcal{E})\geq 1-\delta/2$.
@@
$\frac{\widehat{S}_{t,s}}{\widehat{\sigma}^{2}_{s,\mathrm{con}}}$ $\leq\frac{c_{6}\lambda_{\max}^{s}({\Sigma})h(p,n,s,t)}{c_{4}\lambda_{\max}^{s}({\Sigma})}$ (175)
@@
$C>32(\lambda c_{7}+c_{6})^{2},$ (179)
where $c_{7}$ is defined below and $c_{6}$ is as before.
@@
$\mathcal{E}_{11}$ =$=\left\{\frac{\widehat{\sigma}^{2}_{s,\mathrm{con}}}{\lambda_{\max}^{s}(\Sigma_{1})\wedge\lambda_{\max}^{s}(\Sigma_{2})}\leq c_{7}\right\},$ (180)
$\mathcal{E}_{12}$ $=\bigcap_{i=1,2}\left\{\widehat{\lambda}_{\max}^{s}({\widehat{\Sigma}_{i,t}-\Sigma_{i}})\leq\frac{c_{6}}{2}\lambda_{\max}^{s}({\Sigma_{i}})h(p,n,s,t)\right\},$ (181)
where $c_{7}=1+2c_{0}\log(8/\delta)$ and $c_{6}$ is as before. Define $\mathcal{E}=\mathcal{E}_{11}\cap\mathcal{E}_{12}$. Due to Lemma 8, we have that $\mathbb{P}(\mathcal{E})\geq 1-\delta/2$. On $\mathcal{E}$, we have
@@
$\geq\lambda_{\max}^{s}({\Sigma_{1}-\Sigma_{2}})-c_{6}\left\{\lambda_{\max}^{s}(\Sigma_{1})\vee\lambda_{\max}^{s}(\Sigma_{2})\right\}h(p,n,s,t).$ (183)
@@
$\geq\frac{C}{32}s_{0}^{2}\left\{\log(ep/s_{0})\vee\log\log(8n)\right\}$ (187)
$\geq\frac{C}{32}s^{2}\left\{\log(ep/s)\vee\log\log(8n)\right\}.$ (188)
Since $C\geq 32$, the chain of inequalites preceding Equation (188) imply that
$t\geq s^{2}\left\{\log(ep/s)\vee\log\log(8n)\right\}$, since $\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})$ can never exceed the maximum of $\lambda_{\max}^{s}(\Sigma_{1})$ and $\lambda_{\max}^{s}(\Sigma_{2})$. It follows that
$\lambda_{\max}^{s}({\Sigma_{1}-\Sigma_{2}})$ $\geq\sqrt{\frac{C}{32}}\left\{\lambda_{\max}^{s}(\Sigma_{1})\vee\lambda_{\max}^{s}(\Sigma_{2})\right\}h(p,n,s,t).$ (189)
@@
$\widehat{S}_{t,s}-\lambda\widehat{\sigma}^{2}_{s,\mathrm{con}}h(p,n,s,t)$ $\geq\lambda_{\max}^{s}({\Sigma_{1}-\Sigma_{2}})-c_{6}\left\{\lambda_{\max}^{s}(\Sigma_{1})\vee\lambda_{\max}^{s}(\Sigma_{2})\right\}h(p,n,s,t)$ (190)
$-\lambda c_{7}\left\{\lambda_{\max}^{s}(\Sigma_{1})\vee\lambda_{\max}^{s}(\Sigma_{2})\right\}h(p,n,s,t)$ (191)
$\geq\left(\sqrt{\frac{C}{32}}-c_{6}-\lambda c_{7}\right)\left\{\lambda_{\max}^{s}(\Sigma_{1})\vee\lambda_{\max}^{s}(\Sigma_{2})\right\}h(p,n,s,t)$ (192)
@@
since $C>32(\lambda c_{7}+c_{6})^{2}$. It follows that
@@
6 Supplementary material
In the following, we will for any positive definite matrix $V$ let $\phi_{V}(x)$ denote the density of a multivariate Gaussian with mean zero and covariance matrix $V$. Moreover, we say that a variable $H$ follows a Hypergeometric distribution, denoted $H\sim\text{Hyp}(N,K,n)$, if $H$ has point mass function
$\mathbb{P}(H=i)$ =($=\frac{{K\choose i}{N-K\choose n-i}}{{N\choose n}},$ (196)
for $i=\max(0,n+K-n),\dots,\min(n,K)$.
6.1 Auxiliary Lemmas

### v2

@@ 4.1 A convex relaxation of the k-sparse eigenvalue problem
The testing procedures presented in the previous section have provably strong theoretical performance, but are computationally intractable for all but small values of $p$. Indeed, a key ingredient in these tests is the $s$-sparse eigenvalue of the $p\times p$ matrix $\widehat{\Sigma}_{1,t}-\widehat{\Sigma}_{2,t}$. Seeing as the sparse eigenvalue is NP hard to compute in general (Magdon-Ismail,, 2017), the changepoint testing procedures in the previous section are thus prohibitively computationally costly unless $p$ is very small.
As a remedy to the large computational cost, the $s$-sparse eigenvalue can be approximated by a convex relaxation of the implicit optimization problem via Semidefinite Programming, as is done in Berthet and Rigollet, 2013b for testing for a rank one perturbation in an isotropic covariance matrix. Following Berthet and Rigollet, 2013b and d’Aspremont et al., (2007), we can reformulate the sparse eigenvalue problem as follows. For any symmetric matrix $A\in\mathbb{R}^{p\times p}$, recalling that
@@
The optimization problem in (54) features two sources of non-convexity; the $\ell_{0}$ constraint and the rank constraint. To relax the $\ell_{0}$ constraint, notice that $\left\lVert Z\right\rVert_{1}\leq s$ whenever $\mathrm{Tr}(Z)=1$ and $\left\lVert Z\right\rVert_{0}\leq s^{2}$, since $\left\lVert Z\right\rVert_{1}\leq s\left\lVert Z\right\rVert_{2}=s\mathrm{Tr}(Z)^{1/2}=s\mathrm{Tr}(v^{\top}v)^{1/2}=s$. Thus, a convex relaxation of the $\ell_{0}$ constraint is given by $\left\lVert Z\right\rVert_{1}\leq s$. For the rank constraint, however, there is no obvious convex relaxation. By simply dropping the rank constraint, the optimization problem in (54) can be relaxed to the semidefinite program
@@
where $N(p,s)=\{Z\in\mathbb{R}^{p\times p}\ ;\ Z\succcurlyeq 0,\mathrm{Tr}(Z)=1,\left\lVert Z\right\rVert_{1}\leq s\}$. As a convex problem, it can be solved efficiently using e.g. interior point methods (see Boyd and Vandenberghe,, 2004) or first order methods (see Bach et al.,, 2010), the latter in polynomial time in $p$.
Due to the results in Bach et al., (2010), formalized in Lemma 11 in the Supplementary Material, we have
@@ 4.2 The testing procedure
where the $\widehat{\Sigma}_{i,t}$ for $i=1,2$ are defined in (42). Since we will use (57) to control the behavior of $\widehat{S}_{t,s}$, it suffices to estimate the noise level of $\widehat{S}_{t,s}$ by
$\widehat{\sigma}^{2}_{\mathrm{con}}$ $=\widehat{\lambda}^{1}_{\max}(\widehat{\Sigma}_{1,\lceil\log(ep)\rceil})\wedge\widehat{\lambda}^{1}_{\max}(\widehat{\Sigma}_{2,\lceil\log(ep)\rceil})$ (59)
$=\lVert\widehat{\Sigma}_{1,\lceil\log(ep)\rceil}\rVert_{\infty}\wedge\lVert\widehat{\Sigma}_{2,\lceil\log(ep)\rceil}\rVert_{\infty},$ (60)
where $\widehat{\Sigma}_{i,t}$ is defined in (42) for $i=1,2$ and $t\in[n]$. We remark that $\widehat{\sigma}^{2}_{\mathrm{con}}$ uses the first and last $\lceil\log(ep)\rceil$ observations to estimate the order of the noise level of $\widehat{S}_{t,s}$, which implicitly restricts the sample size to satisfy $n\geq\log(ep)$.
We define the computationally tractable variant of the testing procedure in (49) by
@@
for some $s\in[p]$, where $C>0$ is a constant depending only on $\lambda,\delta,w$ and $u$.
Comparing Proposition 6 to Proposition 5, the computational feasibility of the testing procedure in (61) comes at the price of an increased signal strength required for guaranteed detection of a changepoint. To see this, recall that the computationally unfeasible testing procedure in (43) has a signal strength requirement of order $s\log(ep/s)\vee\log\log(8n)$. In comparison, the computationally feasible variant in (61) has a signal strength requirement of order $s^{2}\{\log(ep)\vee\log\log(8n)\}$. However, when $p\geq\log n$, this discrepancy is likely to be unavoidable for any polynomial-time algorithm. Indeed, in the special case where the changepoint location is known, $\Sigma_{1}=I$ and $\Sigma_{2}=I+\theta vv^{\top}$ for some $\theta>0$ and $v\in S^{p-1}_{s}$, the covariance changepoint problem reduces to a sparse PCA problem. For this problem, Berthet and Rigollet, 2013a showed that if (64) were to be improved by a randomized polynomial-time algorithm, then such an algorithm could be modified to detect a Planted Clique in a regime widely conjectured to be computationally infeasible (see also Brennan and Bresler,, 2019). For the case where $n>e^{p}$, however, it is unclear whether the rate $s^{2}\log\log(8n)$ in (64) is improvable for polynomial-time algorithms.
@@ 5.1 Proofs of results from Section 2
where we in the second inequality used that $\log\log(8n)\geq 1$, and last inequality used that $c_{1}C^{1/2}/(2c_{3})-\lambda>0$ and $t\leq(16c_{0}^{2}\vee 1)\log(8/\delta)$. Due to the definition of $C$, the right-hand side of (91) is positive, and therefore $\psi_{\lambda}(X)=1$ holds on the event $\mathcal{E}$ also when $t\leq(16c_{0}^{2}\vee 1)\log(8/\delta)$.
@@
The proof of Proposition 2 adopts techniques from Liu et al., (2021). Fix any $\sigma>0,w\geq(2\pi)^{-1/2}$ and $u>0$ sufficiently large. To prove Proposition 2 we will impose a prior on the alternative hypothesis parameter space and bound the total variation distance between the null distribution and the mixture distribution induced by the prior. Since $w\geq(2\pi)^{-1/2}$ and $u$ is sufficiently large, we can take the distribution of the $X_{i}$ to be Gaussian. Given any $0\leq\eta\leq 1$, it then suffices by Lemma 13 to find an absolute constant $c>0$ and a probability measure $\nu$ with $\text{supp}(\nu)\subseteq\Theta(n,c\log\log(8n))$ such that $\chi^{2}(f_{1},f_{0})\leq\eta,$ where $f_{0}$ and $f_{1}$ are the densities of $X\sim\text{N}_{n}(0,V)$ when $V=\sigma^{2}I$ and $V\sim\nu$, respectively.
@@
    Let $\Delta=2^{l}$, where $l$ is sampled uniformly from the set
    $\left\{0,1,2,\ldots,{\lfloor\log_{2}(n/2)\rfloor\}}\right\}$.
@@
$\min(\Delta,n-\Delta)\left\{\left(\frac{\left|\sigma_{1}^{2}-\sigma_{2}^{2}\right|}{\sigma_{1}^{2}\wedge\sigma^{2}_{2}}\right)\wedge\left(\frac{\left|\sigma_{1}^{2}-\sigma_{2}^{2}\right|}{\sigma_{1}^{2}\wedge\sigma^{2}_{2}}\right)^{2}\right\}=\rho,$ (94)
@@
where $l_{i}=\log_{2}(\Delta_{i})$ for $i=1,2$. We bound the two terms on the right-hand side of (100) separately. For the first term, we have
@@
Now choose $c\leq\log(16/\eta)^{-1}/2\wedge 1$. Then if $\log\log(8n)<c^{-1}$, the right-hand side of (102) is zero since $c\log\log(8n)<1$. If we instead have $\log\log(8n)\geq c^{-1}$, the right-hand size of (102) is bounded above by
@@
Now we bound the second term in the right-hand side of (100). We have
@@
Note that $l_{1},l_{2}$ are sampled independently and uniformly from an integer interval with cardinality $a(n)=\left\lfloor\log_{2}(n/2)\right\rfloor+1$, from which it follows that $\mathbb{P}(|l_{1}-l_{2}|=x)\leq 2a(n)^{-1}$ for any $x\in\mathbb{N}\cup\{0\}$. We claim that the first term at the right-hand side of (108) is bounded above by
@@
$\mathbb{E}_{l_{1},l_{2}}\exp\left\{\frac{c}{2}\log\log(8n)\right\}\mathbbm{1}\left\{\left|l_{1}-l_{2}\right|=0\right\}$ (116)
$=$ $\exp\left\{\frac{c}{2}\log\log(8n)\right\}\mathbb{P}_{l_{1},l_{2}}(|l_{1}-l_{2}|=0)$ (117)
@@
For the second term at the right-hand side of (108), using that $c\leq 1$, we have
@@
Now let $b_{\eta}=\ \underset{n\geq 2}{\sup}\ \frac{\log\log(8n)}{\log(8n)^{\eta\log(2)/18}}<\infty$. Then for the second term on the right-hand side of (108), we have
@@ 5.2 Proofs of results from Section 3
    Let $\Delta=2^{l}$, where $l$ is sampled uniformly from the set
    $\left\{0,1,2,\ldots,\lfloor\log_{2}(n/2)\rfloor\right\}.$
@@
$\min(t_{0},n-t_{0})\left\{\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)\wedge\left(\frac{\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}{\sigma^{2}-\left\lVert\Sigma_{1}-\Sigma_{2}\right\rVert_{\mathrm{op}}}\right)^{2}\right\}=\rho,$ (129)
@@
Using similar arguments as above to bound the two terms at right-hand side of Equation (134), the event
@@
$\mathcal{E}_{5}$ $=\bigcap_{\begin{subarray}{c}s\in\mathcal{S};\\ \gamma(s)\leq n\end{subarray}}\left\{c_{4}\leq\frac{\widehat{\sigma}^{2}_{s}}{\lambda_{\max}^{s}({\Sigma})}\right\},$ (153)
@@
Without loss of generality, we may assume that $\min(t_{0},n-t_{0})=t_{0}$. Note that there exists some $t\in\mathcal{T}$ such that $t_{0}/2\leq t\leq t_{0}$ and some $s\in\mathcal{S}$ such that $s_{0}/2\leq s\leq s_{0}$ and $s\log(ep/s)\leq n$. Fixing this $s$ and $t$, we define the events
@@
Since $C\geq 32$, the chain of inequalities preceding Equation (168) imply that $t\geq s\log(ep/s)\vee\log\log(8n)$, since $\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})\leq\lambda_{\max}^{s}(\Sigma_{1})\vee\lambda_{\max}^{s}(\Sigma_{2})$, due to Lemma 12. It follows that $r(p,n,s,t)=t^{-1/2}\{s\log(ep/s)\vee\log\log(8n)\}^{1/2}$ and thus
@@ 5.3 Proofs of results from Section 4
$\mathcal{E}_{9}$ =$=\left\{c_{4}\leq\frac{\widehat{\sigma}^{2}_{\mathrm{con}}}{\lambda_{\max}^{1}({\Sigma})}\right\},$ (176)
$\mathcal{E}_{10}$ $=\bigcap_{t\in\mathcal{T}}\bigcap_{s\in\mathcal{S}}\left\{\widehat{\lambda}_{\max}^{s}({\widehat{\Sigma}_{1,t}-\widehat{\Sigma}_{2,t}})\leq c_{6}\lambda_{\max}^{1}({\Sigma})h(p,n,s,t)\right\},$ (177)
where $\widehat{\sigma}^{2}_{\mathrm{con}}$ is given in (60), $\widehat{\Sigma}_{i,t}$ is given in (42), $\mathcal{T}$ is given in (21), $\mathcal{S}$ is given in (50), $h(p,n,s,t)$ is given in (62), $c_{4}=\delta^{2}(\delta^{2}+16\delta+64)^{-1}(2e\pi w^{2})^{-1}$, $c_{6}=c_{0}\{1+\log(4/\delta)\}$, and $c_{0}$ is the constant from Lemma 2 depending only on $u>0$. Define the event $\mathcal{E}=\mathcal{E}_{9}\cap\mathcal{E}_{10}$. By Lemma 7 we have $\mathbb{P}(\mathcal{E})\geq 1-\delta/2$.
@@
$\frac{\widehat{S}_{t,s}}{\widehat{\sigma}^{2}_{\mathrm{con}}}$ $\leq\frac{c_{6}\lambda_{\max}^{1}({\Sigma})h(p,n,s,t)}{c_{4}\lambda_{\max}^{1}({\Sigma})}$ (178)
@@
$C>32(\lambda c_{3}+c_{6})^{2},$ (182)
where $c_{3}=1+c_{0}\log(1+4/\delta)$ and $c_{6}$ is as before.
@@
$\mathcal{E}_{11}$ =$=\left\{\frac{\widehat{\sigma}^{2}_{\mathrm{con}}}{\lambda_{\max}^{1}(\Sigma_{1})\wedge\lambda_{\max}^{1}(\Sigma_{2})}\leq c_{3}\right\},$ (183)
$\mathcal{E}_{12}$ $=\bigcap_{i=1,2}\left\{\widehat{\lambda}_{\max}^{s}({\widehat{\Sigma}_{i,t}-\Sigma_{i}})\leq\frac{c_{6}}{2}\lambda_{\max}^{1}({\Sigma_{i}})h(p,n,s,t)\right\},$ (184)
where $c_{3}$ and $c_{6}$ are as before. Define $\mathcal{E}=\mathcal{E}_{11}\cap\mathcal{E}_{12}$. Due to Lemma 8, we have that $\mathbb{P}(\mathcal{E})\geq 1-\delta/2$. On $\mathcal{E}$, we have
@@
$\geq\lambda_{\max}^{s}({\Sigma_{1}-\Sigma_{2}})-c_{6}\left\{\lambda_{\max}^{1}(\Sigma_{1})\vee\lambda_{\max}^{1}(\Sigma_{2})\right\}h(p,n,s,t).$ (186)
@@
$t\left(\frac{\lambda_{\max}^{s}({\Sigma_{1}-\Sigma_{2}})}{\lambda_{\max}^{1}(\Sigma_{1})\vee\lambda_{\max}^{1}(\Sigma_{2})}\right)^{2}$ $\geq\frac{t}{16}\left(\frac{\lambda_{\max}^{s_{0}}({\Sigma_{1}-\Sigma_{2}})}{\lambda_{\max}^{s_{0}}(\Sigma_{1})\vee\lambda_{\max}^{s_{0}}(\Sigma_{2})}\right)^{2}$ (188)
@@
$\geq\frac{C}{32}s_{0}^{2}\left\{\log(ep)\vee\log\log(8n)\right\}$ (190)
$\geq\frac{C}{32}s^{2}\left\{\log(ep)\vee\log\log(8n)\right\}.$ (191)
Since $C\geq 32$, the chain of inequalities preceding Equation (191) imply that
$t\geq\left\{\log(ep)\vee\log\log(8n)\right\}$, since $\lambda_{\max}^{s}(\Sigma_{1}-\Sigma_{2})$ can never exceed the maximum of $\lambda_{\max}^{s}(\Sigma_{1})$ and $\lambda_{\max}^{s}(\Sigma_{2})$. It follows that $h(p,n,s,t)=st^{-1/2}\{\log(ep)\vee\log\log(8n)\}^{1/2}$, and thus
$\lambda_{\max}^{s}({\Sigma_{1}-\Sigma_{2}})$ $\geq\sqrt{\frac{C}{32}}\left\{\lambda_{\max}^{1}(\Sigma_{1})\vee\lambda_{\max}^{1}(\Sigma_{2})\right\}h(p,n,s,t).$ (192)
@@
$\widehat{S}_{t,s}-\lambda\widehat{\sigma}^{2}_{\mathrm{con}}h(p,n,s,t)$ (193)
$\geq$ $\lambda_{\max}^{s}({\Sigma_{1}-\Sigma_{2}})-c_{6}\left\{\lambda_{\max}^{1}(\Sigma_{1})\vee\lambda_{\max}^{1}(\Sigma_{2})\right\}h(p,n,s,t)$ (194)
-- $\lambda c_{3}\left\{\lambda_{\max}^{1}(\Sigma_{1})\vee\lambda_{\max}^{1}(\Sigma_{2})\right\}h(p,n,s,t)$ (195)
$\geq$ $\left(\sqrt{\frac{C}{32}}-c_{6}-\lambda c_{3}\right)\left\{\lambda_{\max}^{1}(\Sigma_{1})\vee\lambda_{\max}^{1}(\Sigma_{2})\right\}h(p,n,s,t)$ (196)
@@
since $C>32(\lambda c_{3}+c_{6})^{2}$. It follows that
@@
6 Auxiliary Lemmas